OpenStudy (anonymous):

Write this in an equation in standard form please? (5,1), (-3, 4)

OpenStudy (anonymous):

5.1 -3.5 are you doing intigers?

OpenStudy (anonymous):

Do you mean the y=mx+b Equationg?

OpenStudy (anonymous):

*equation

OpenStudy (anonymous):

If so, use the (y2-y1) over (x2-x1) to find the slope, then just substitute the values into the equation.

OpenStudy (anonymous):

This is what i have to do: Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. And i already have the slope-intercept and point slope. i dont know how to get syandard form?

OpenStudy (triciaal):

equation in standard form Ax + By = C or Ax + By + C = 0 you have 2 points (x1, y1) and (x2, y2) slope, m = change in y / corresponding change in x = (y2 - y1)/(x2 - x1) using one of the points given and y = mx + b where b is the intercept find b slope-intercept format now write the equation for any point (x, y) rearrange to put in standard form note that you could also use point slope format to write the equation (y-y1) = m(x-x1)

OpenStudy (anonymous):

woa, okay. thanks! but im still confused?

OpenStudy (anonymous):

okay, wait, would it be 5x+1y=15?

OpenStudy (triciaal):

I have to actually do it what is the slope?

OpenStudy (anonymous):

lets see, i think -3/8?

OpenStudy (triciaal):

(5,1), (-3, 4) 4-1/(-3 - 5) = 3/-8 y - -3 = -3/8(x - 4) y + 3 = -3/8 x + 12

OpenStudy (triciaal):

3x + 8y = 9

OpenStudy (anonymous):

So, thats the standard form? and thank you so much!

OpenStudy (triciaal):

I've learnt that some books use Ax + By = C and some will use Ax + By + C= 0

OpenStudy (anonymous):

Okay

OpenStudy (triciaal):

here I used point-slope to use slope intercept you would choose a given point example (5, 1) with y = mx + b solve for b then generalize do you have any questions about the steps?

OpenStudy (anonymous):

Okay, i tried finding the slope-intercept and point slope- can you check and make sure its correct? Slope intercept form- y =-3/8x+3 and point slope form y-1=-3/8(x-5)

OpenStudy (triciaal):

using Slope intercept form- y =-3/8x+3 y = mx + b 5 = -3/8(1) + b b = 5 + 3/8 = 43/8 then substitute using (x, y) y = -3/8(x ) + 43/8 multiply everything by 8 to get rid of fraction 8y = -3x + 43 3x + 8y = 43

OpenStudy (triciaal):

error above (5,1), (-3, 4) 4-1/(-3 - 5) = 3/-8 y - -3 = -3/8(x - 4) y + 3 = -3/8 x + 12 should be 12/8 = 3/2

OpenStudy (anonymous):

so, 12/8 = 3/2 is the slope-intercept equation?

OpenStudy (triciaal):

3x + 8y = 9 with correction (3/2 + 3)*8

OpenStudy (anonymous):

Okay. so then i solve the rest? would it be Point slope: m= 1 - 4 (5+3) Standard: 5x + 1y = 15 Slope intercept: 1 = 1 1/2

OpenStudy (triciaal):

3x + 8y = found BIG ERROR (5,1) x = 5 y = 1 sorry

OpenStudy (anonymous):

Okay, so now im confused? haha

OpenStudy (triciaal):

start right here slope intercept form y = m x + b y = 1 x = 5 m = -3/8 1 = -3/8 (5) + b solve for b b = 1 + 15/8 = 23/8 using any (x, y) y = -3/8(x) + 23/8 multiply by 8 (to get rid of fraction ) then rearrange to standard form 3x + 8y = 23

OpenStudy (triciaal):

point-slope y - 1 = -3/8(x -5) y - 1 = -3/8 x + 15/8 multiply by 8 and rearrange 3x + 8y = (15 +8) = 23

OpenStudy (anonymous):

I have to go, but ill get back on later. Thank you for all your help!

OpenStudy (triciaal):

so, now you see 1) the importance of using the correct x and y 2) you can use either point-slope or slope intercept 3) how to get equation of line and rearrange to standard form

OpenStudy (triciaal):

sorry it took so long you are always welcome